OpenStudy (anonymous):

What is the slope of the line passing through points A and B? Point A is plotted at (2) and point B is plotted at (3,3). Please explain on how i would do this! A. -3 B. -1/5 C. 1/4 D. 1/3

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

Points A: (x1,y1) Points B: (x2,y2)

OpenStudy (anonymous):

I dont understand?

OpenStudy (anonymous):

Ok, you know coordinates have (x,y) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, the slope is trying to find the rise/run

OpenStudy (anonymous):

ummmmmm okay...

OpenStudy (anonymous):

So you take your y-coordinate from B - the y-coordinate from A

OpenStudy (anonymous):

over the x-coordinate from B - the x-coordinate from A

OpenStudy (anonymous):

Im still not getting this

OpenStudy (anonymous):

i got 1 for each of them

OpenStudy (anonymous):

SInce the A coordinates are (2,0) and the B is (3,3) 3-0/3-2 = 3/1 = 3

OpenStudy (anonymous):

Mhmm, did you forget a coordinate in your question?

OpenStudy (anonymous):

for Point A

OpenStudy (anonymous):

wait so its -3?

OpenStudy (anonymous):

Well your coordinates for point A is (2). Makes no sense.

OpenStudy (anonymous):

Unless you mean (2,0) which is what I assumed.

OpenStudy (anonymous):

ya becasue you know where the zeros at in a graph well you go up 2 and thats it

OpenStudy (anonymous):

thats why i just puit 2

OpenStudy (anonymous):

Oh! Then it's (0,2). Never leave it alone when they're coordinates.

OpenStudy (anonymous):

Then it's 3-2/ 3-0 = 1/3

OpenStudy (anonymous):

OOOOOOOOOOOOOOOOOOOOOOOOOOOOOO

OpenStudy (anonymous):

I get it!!!